# Changes available at 2026-10-15 11:30
C-01: Nia is unavailable 12:30-13:30. Today's AV capacity is now 30 minutes. Nobody accepted substitute cover. Existing deadlines have not moved.
C-02: Uma clarified SR-208: panel is Friday 2026-10-16 14:00. Put it in tomorrow's queue; no tomorrow capacity has been supplied.
C-03: New request SR-211, source R-211, received 11:20 from Vera: damaged power strip in Room Birch, needed by 13:00 today, Facilities assessment estimate 30 minutes. Reported sparking; request isolation/notification via site safety contact; no confirmation of completed isolation. Do not energize or repair.
C-04: Facilities capacity remains 90 minutes. Do not add overtime. Arun must assess SR-211 and SR-203 before ordinary room reset SR-206.
C-05: Dev has not approved SR-204. SR-210 still lacks a sponsor. IT has not acknowledged SR-205. The cable remains an additional item; no cancellations were received.
C-06: Lila must escalate displaced microphone SR-207 and room reset SR-206 to requesters for alternatives before their needed-by time. This is a follow-up duty, not confirmed alternative fulfilment.
